# Enumeration of maps with tight boundaries and the Zhukovsky transformation

### v1

@@ 2.2 Enumerative consequences
For $g\geq 0$ and $L_{1},\ldots,L_{n}$ positive integers ($n\geq 1$), let us denote by $F^{(g)}_{L_{1},\ldots,L_{n}}$ the generating function of maps of genus $g$ with $n$ rooted boundary-faces of lengths $L_{1},\ldots,L_{n}$, and by $\hat{T}^{(g)}_{L_{1},\ldots,L_{n}}$ the generating function of those maps whose boundaries are tight. Then, for $(g,n)\neq(0,1)$ or equivalently $\chi=2-2g-n\leq 0$, we have
@@
The usefulness of this statement comes from the fact that $A_{L,\ell}$ admits an explicit expression.
@@
[irredmaps, Section 9.3] For any two positive integers $L,\ell$, we have
@@
Note that the notation from [irredmaps] differs from that of the present paper: precisely, Proposition 2.4 is implied by formula (9.19) in this reference, in the case $d=0$, $d^{\prime}=\ell$ and $n=L$.
@@
which is a formal power series in $x_{1}^{-1},\ldots,x_{n}^{-1}$ ($n\geq 1$). Then, following [Eynard2016, Definition 3.3.1, p. 87], we define for $(g,n)\neq(0,1),(0,2)$ [^2] Some conventional terms have to be added in these cases., or equivalently $\chi=2-2g-n<0$, the quantity
@@
can be seen as a formal power series in $z^{-1}$ without constant coefficient, so that for any $\ell$, $x^{\prime}(z)/x(z)^{\ell+1}$ is a series containing only powers of $z^{-1}$ larger than or equal to $\ell+1$. This shows that
@@
$[x^{-1}]F(x)=[z^{-1}]F(x(z))x^{\prime}(z).$ (14)
Indeed, by linearity it suffices to check this identity for $F(x)=x^{k}$, $k\in\mathbb{Z}$: both sides are equal to $1$ for $k=-1$, and vanish for $k\neq-1$ (since $x(z)^{k}x^{\prime}(z)=\frac{d}{dz}\frac{x(z)^{k+1}}{k+1}$ and a derivative contains no monomial in $z^{-1}$). Applying the identity for each variable of the multivariate formal Laurent series $x_{1}^{L_{1}}\cdots x_{n}^{L_{n}}W_{n}^{(g)}(x_{1},\ldots,x_{n})$, we obtain
@@
where $A_{L,\ell}$ is as in (6). Comparing with (5), and noting that the matrix $(A_{L,\ell})_{L,\ell\geq 1}$ is unitriangular hence invertible, we get the wanted equality (11). ∎
@@ 2.3 Allowing boundary-vertices
We now consider the enumerative consequences of Theorem 2.7. In order to put boundary-faces and boundary-vertices on a same footing, we denote by $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ the generating function of maps of genus $g$ with $n$ unrooted tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$. It is related to the generating function $\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ considered above by
@@
as there are $\ell_{1}\cdots\ell_{n}$ ways to root the boundaries, up to the symmetry factor corresponding to the inverse of the automorphism group order [^3] Note that the objects enumerated by the generating series $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ are not rooted, and therefore can have nontrivial automorphism groups. This is why we have to weigh them by the inverse of the automorphism group order, as discussed in the introduction.. In particular, there are zero ways if one $\ell_{i}$ vanishes: as said in Section 1.2, a boundary-vertex cannot be rooted. Thus, $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ contains a priori more information than $\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. In particular, for given integers $m,s\geq 0$ not both equal to 00, and positive integers $\ell_{1},\ldots,\ell_{m}$, the generating function of maps with $m$ rooted tight boundary-faces of lengths $\ell_{1},\ldots,\ell_{m}$ and with $s$ boundary-vertices is given by $(\ell_{1}\cdots\ell_{m})T^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}$, with $s$ zero indices. With this remark at hand, we can now state the generating function counterpart to Theorem 2.7 as follows.
@@
where there are $s$ indices equal to 00 in the last term, and where we reuse the notations from Corollary 2.3. For $m=0$, the relation reads
@@
with $s$ indices equal to 00, and where $F^{(g)}$ is the generating function of maps of genus $g$ without boundaries.
@@
For $m=0$, we have $\omega^{(g)}_{0,s}=W_{0,s}^{(g)}={\partial^{s}F^{(g)}}/{\partial t^{s}}$. Corollary 2.8 admits the following consequence, analog to Theorem 2.5, whose proof is adapted from the latter in a straightforward way.
@@
Theorems 2.5 and 2.9 provide combinatorial interpretations for the coefficients of $\omega^{(g)}_{n}$ and $\omega^{(g)}_{m,s}$ in terms of generating series of maps with tight boundaries. However, they do not provide any information about the structural properties of $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. This is the goal of the upcoming Section 4, where we will see that these are quasi-polynomials in the variables $\ell_{1},\ldots,\ell_{n}$, and that $\tau^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}$ (with $s$ zero indices) indeed corresponds to the evaluation of this quasi-polynomial function of $n=m+s$ variables when the last $s$ ones are set to 00.
@@ 2.4 A formula for the series of planar bipartite maps with tight boundaries
Our input is the Collet-Fusy formula [CoFu12, Theorem 1.1] for the generating function of planar bipartite maps with $n$ rooted, non necessarily tight, boundaries of prescribed lengths. In our present notations, it reads
@@
for any $n\geq 1$ and positive integers $L_{1},\ldots,L_{n}$. Note that the formula makes sense for $n=1$, upon understanding that $\frac{\partial^{-1}}{\partial t^{-1}}$ means integrating over $t$.
We will deduce from the Collet-Fusy formula an expression for $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\rvert_{\mathrm{bip}}$, using Corollary 2.8. Interestingly, our expression involves a certain family of polynomials that we encountered previously in [polytightmaps]. Namely, for any integers $k\geq 0$ and $n\geq 1$, let us consider the multivariate polynomial in the variables $\ell_{1},\ldots,\ell_{n}$
@@
where, in the right-hand side, all factors except the first one are $q_{k_{i}}$’s, and where the univariate polynomials $p_{k}$ and $q_{k}$ are defined by
@@
with the convention $p_{0}(\ell)=q_{0}(\ell)=1$, and with $\binom{x}{k}=x(x-1)\cdots(x-k+1)/k!$ viewed as a polynomial in $x$. Clearly, $p_{k}(\ell_{1},\ell_{2},\ldots,\ell_{n})$ is a polynomial in the variables $\ell_{1}^{2},\ldots,\ell_{n}^{2}$. Moreover, as discussed in [polytightmaps, Proposition 2.1], it is a symmetric polynomial and it satisfies the consistency relation
@@
Combinatorially, $B_{n,k}(r_{1},r_{2},\ldots)$ is the generating series of partitions of an $n$-element set into $k$ blocks, where we attach a weight $r_{i}$ to eack block of size $i$. Table 1 lists the first few Bell polynomials. They appear in the following classical formula:
@@
See for instance the discussion in [FlSe09, Example III.24]. We may now state the main result of this section:
@@
For $n=3$ we recover $T^{(0)}_{2\ell_{1},2\ell_{2},2\ell_{3}}\rvert_{\mathrm{bip}}=R^{\ell_{1}+\ell_{2}+\ell_{3}-1}R^{\prime}-t^{-1}\delta_{\ell_{1}+\ell_{2}+\ell_{3},0}$ consistently with [triskell, Theorem 1.1], while for $n=4$ we find
@@
When setting the weights $t_{2},t_{4},\ldots$ for inner faces to zero, we have $R=t$, $R^{\prime}=1$ and $R^{(j)}=0$ for $j\geq 2$. Thus, in this case we have $b_{n,k}=t^{-k}\delta_{n,k}$, hence $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}$ reduces to $(n-3)!p_{n-3}(\ell_{1},\ldots,\ell_{n})t^{\ell_{1}+\ldots+\ell_{n}-n+2}$ if at least one $\ell_{i}$ is non-zero, and vanishes otherwise since $p_{n-3}(0,\ldots,0)=(-1)^{n-3}$ so the rightmost term in (32) cancels the result. This is consistent with [polytightmaps, Theorem 2.3] since the tight maps considered in this reference are nothing but maps with tight boundaries and without inner faces. Note that the exponent of $t$ is consistent with Euler’s relation.
@@
Let us first consider the case where all $\ell_{i}$ are zero: then $T^{(0)}_{0,\ldots,0}$ is simply the generating function of bipartite planar maps with $n\geq 3$ marked vertices. For $n=3$ it is equal to $\frac{\partial}{\partial t}\ln(R/t)$, see e.g. [triskell, Appendix A], and for larger $n$ we simply have to take more derivatives with respect to $t$, namely
@@
Let us now assume that at least one $\ell_{i}$ is non-zero, and let $m\geq 1$ be the number of non-zero $\ell_{i}$, and $s=n-m$ the number of zero $\ell_{i}$. Without loss of generality, we can assume that $\ell_{1},\ldots,\ell_{m}>0$ and $\ell_{i}=0$ for $i>m$. We note that, in the bipartite case, the trumpet generating function of Proposition 2.4 vanishes whenever its indices are of different parities (since $S=0$), and reads for even indices
@@
where $(L)_{k}:=L(L-1)\cdots(L-k+1)$ denotes the falling factorial. We will rewrite this expression in the same form as the right-hand side of (38), so as to identify $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{m},0,\ldots,0}\right|_{\mathrm{bip}}$. To this end, we proceed as in [polytightmaps, Section 3.1] and use the Chu-Vandermonde identity to write
@@
Now, by [polytightmaps, Lemma 3.1] and by (37) [^4] Beware that the notation $A_{\ell,m}$ used in [polytightmaps] differs from that of the present paper., we have for $L>0$
@@
and by (28), we may replace $p_{k-1}(\ell_{1},\ldots,\ell_{m})$ by $p_{k-1}(\ell_{1},\ldots,\ell_{m},0,\ldots,0)$, in order to have $n$ variables. Comparing this expression with (38), using the invertibility of the unitriangular matrix $(A_{2L,2\ell})_{L,\ell\geq 1}$ and doing a change of variable $k-1\to k$ finally gives the wanted formula (32), since the term $\frac{(-1)^{n}(n-3)!}{t^{n-2}}$ is absent for $s\geq 1$. ∎
A variant of the Collet-Fusy formula (25) holds in the quasi-bipartite case where exactly two of the boundaries have odd lengths, i.e. when exactly two of the $L_{i}$, say $L_{1}$ and $L_{2}$, are half-integers. In this case, the factors $\binom{2L_{1}}{L_{1}}\binom{2L_{2}}{L_{2}}$ appearing in the right-hand side should be replaced by $4\binom{2L_{1}-1}{L_{1}-1/2}\binom{2L_{2}-1}{L_{2}-1/2}$. Correspondingly, we have the following analogue of Theorem 2.11:
@@
where, following [polytightmaps, Section 2.1.3], we define
@@
Starting with the quasi-bipartite variant of the Collet-Fusy formula, we proceed as in the proof of Theorem 2.11. Using Faà di Bruno’s formula, we obtain the expression (39) with the binomial factors $\binom{2L_{1}}{L_{1}}\binom{2L_{2}}{L_{2}}$ modified as discussed above. Now, we expand the falling factorial as
@@
which leads to a variant of (41) in which the factors $\binom{L_{1}-1}{k_{1}}\binom{2L_{1}}{L_{1}}\binom{L_{2}}{k_{2}}\binom{2L_{2}}{L_{2}}$ are replaced by $4\binom{L_{1}-1/2}{k_{1}}\binom{2L_{1}-1}{L_{1}-1/2}\binom{L_{2}-1/2}{k_{2}}\binom{2L_{2}-1}{L_{2}-1/2}$. By [polytightmaps, Equation (3.9)], we have for $L$ a positive half-integer
@@
Let us consider the case of four boundaries. We have $\tilde{p}_{1}(\ell_{1},\ell_{2};\ell_{3},\ell_{4})=\ell_{1}^{2}+\cdots+\ell_{4}^{2}-\frac{1}{2}$ and $\tilde{p}_{0}(\ell_{1},\ell_{2};\ell_{3},\ell_{4})=1$, hence from Proposition 2.13 we see that, in the quasi-bipartite case, the term $-1$ appearing in the expression (34) should be replaced by $-\frac{1}{2}$. In contrast, we found after a lengthy computation (which we omit here) that (34) holds as is when all $\ell_{i}$ are half-integers, i.e. when all four boundaries have odd lengths. This is consistent with the discussion of the lattice count polynomial $N_{0,4}$ at the end of [Norbury2010, Section 1], since $T^{(0)}_{2\ell_{1},2\ell_{2},2\ell_{3},2\ell_{4}}$ reduces to $N_{0,4}(2\ell_{1},2\ell_{2},2\ell_{3},2\ell_{4})$ when we set the weights for inner faces to zero.

### v2

@@ 2.2 Enumerative consequences
For $g\geq 0$ and $L_{1},\ldots,L_{n}$ positive integers ($n\geq 1$), let us denote by $F^{(g)}_{L_{1},\ldots,L_{n}}$ the generating function of maps of genus $g$ with $n$ rooted boundary-faces of lengths $L_{1},\ldots,L_{n}$, and by $\hat{T}^{(g)}_{L_{1},\ldots,L_{n}}$ the generating function of those maps whose boundaries are tight [^2] By convention we use a hat symbol to indicate that the boundaries are rooted. Later on we will consider the hatless generating function $T^{(g)}_{L_{1},\ldots,L_{n}}$ where the boundaries are unrooted, see (17).. Then, for $(g,n)\neq(0,1)$ or equivalently $\chi=2-2g-n\leq 0$, we have
@@
Note that, by their very definition, $F^{(g)}_{L_{1},\ldots,L_{n}}$ and $\hat{T}^{(g)}_{L_{1},\ldots,L_{n}}$ are symmetric functions of $L_{1},\ldots,L_{n}$. The usefulness of the statement above comes from the fact that $A_{L,\ell}$ admits an explicit expression:
@@
[BG14, Section 9.3] For any two positive integers $L,\ell$, we have
@@
Note that the notation from [BG14] differs from that of the present paper: precisely, Proposition 2.4 is implied by formula (9.18) in this reference, in the case $d=0$, $d^{\prime}=\ell$ and $n=L$. The right-hand side of (6) is then nothing but the three-step path generating function $P_{\ell}(L;R^{(0)},S^{(0)})$, beware that the variable $z$ used here does not have the same meaning as in [BG14].
@@
which is a formal power series in $x_{1}^{-1},\ldots,x_{n}^{-1}$ ($n\geq 1$). Then, following [Eyn16, Definition 3.3.1, p. 87], we define for $(g,n)\neq(0,1),(0,2)$ [^3] Some conventional terms have to be added in these cases., or equivalently $\chi=2-2g-n<0$, the quantity
@@
can be seen as a formal power series in $z^{-1}$ without constant coefficient, so that for any $L$, $x^{\prime}(z)/x(z)^{L+1}$ is a series containing only powers of $z^{-1}$ larger than or equal to $L+1$. This shows that
@@
$x^{-1}]F(x)=[z^{-1}]F(x(z))x^{\prime}(z).$ (14)
Indeed, by linearity it suffices to check this identity for $F(x)=x^{k}$, $k\in\mathbb{Z}$: both sides are equal to $1$ for k=1k=-1, and vanish for $k\neq-1$ (since $x(z)^{k}x^{\prime}(z)=\frac{d}{dz}\frac{x(z)^{k+1}}{k+1}$ and a derivative contains no monomial in $z^{-1}$). Applying the identity for each variable of the multivariate formal Laurent series $x_{1}^{L_{1}}\cdots x_{n}^{L_{n}}W_{n}^{(g)}(x_{1},\ldots,x_{n})$, we obtain
@@
where $A_{L,\ell}$ is as in (6). Comparing with (5), and noting that the matrix $(A_{L,\ell})_{L,\ell\geq 1}$ is unitriangular hence invertible, we get equality (11), as desired. ∎
@@ 2.3 Allowing boundary-vertices
We now consider the enumerative consequences of Theorem 2.7. In order to put boundary-faces and boundary-vertices on a same footing, we denote by $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ the generating function of maps of genus $g$ with $n$ unrooted tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$. As the generating function $\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ considered above, it is symmetric in $\ell_{1},\ldots,\ell_{n}$, and these two quantities are related by
@@
as there are $\ell_{1}\cdots\ell_{n}$ ways to root the boundaries, up to the symmetry factor corresponding to the inverse of the automorphism group order [^4] Note that the objects enumerated by the generating series $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ are not rooted, and therefore can have nontrivial automorphism groups. This is why we have to weigh them by the inverse of the automorphism group order, as discussed in the introduction.. In particular, there are zero ways if one $\ell_{i}$ vanishes: as said in Section 1.2, a boundary-vertex cannot be rooted. Thus, $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ contains a priori more information than $\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. In particular, for given integers $m,s\geq 0$ not both equal to 0, and positive integers $\ell_{1},\ldots,\ell_{m}$, the generating function of maps with $m$ rooted tight boundary-faces of lengths $\ell_{1},\ldots,\ell_{m}$ and with $s$ boundary-vertices is given by $(\ell_{1}\cdots\ell_{m})T^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}$, with $s$ zero indices. With this remark at hand, we can now state the generating function counterpart to Theorem 2.7 as follows.
@@
where there are $s$ indices equal to 0 in the last term, and where we reuse the notations from Corollary 2.3. For $m=0$, the relation reads
@@
with $s$ indices equal to 0, and where $F^{(g)}$ is the generating function of maps of genus $g$ without boundaries.
@@
For $m=0$, we have $\omega^{(g)}_{0,s}=W_{0,s}^{(g)}={\partial^{s}F^{(g)}}/{\partial t^{s}}$. Corollary 2.8 admits the following consequence, analogous to Theorem 2.5, whose proof is adapted from the latter in a straightforward way.
@@
Theorems 2.5 and 2.9 provide combinatorial interpretations for the coefficients of $\omega^{(g)}_{n}$ and $\omega^{(g)}_{m,s}$ in terms of generating series of maps with tight boundaries. However, they do not provide any information about the structural properties of $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. This is the goal of the upcoming Section 4, where we will see that, for 22gn<02-2g-n<0, these are quasi-polynomials in the variables $\ell_{1},\ldots,\ell_{n}$, and that $\tau^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}$ (with $s$ zero indices) indeed corresponds to the evaluation of this quasi-polynomial function of $n=m+s$ variables when the last $s$ ones are set to 0.
Remark 2.10.
The case $(g,n)=(0,2)$ is special, since we have for $\ell_{1}>0$
$\tau^{(0)}_{\ell_{1},\ell_{2}}=\frac{\delta_{\ell_{1},\ell_{2}}}{\ell_{1}},\qquad T^{(0)}_{\ell_{1},\ell_{2}}=\frac{R^{\ell_{1}}\delta_{\ell_{1},\ell_{2}}}{\ell_{1}},\qquad\hat{T}^{(0)}_{\ell_{1},\ell_{2}}=\ell_{1}R^{\ell_{1}}\delta_{\ell_{1},\ell_{2}}$ (25)
which are not quasi-polynomials in $\ell_{1},\ell_{2}$. Indeed, it is clear that these quantities vanish for $\ell_{1}\neq\ell_{2}$, and we get $\ell_{1}T^{(0)}_{\ell_{1},\ell_{1}}=R^{\ell_{1}}$ by taking $d=0$, $n=d^{\prime}=\ell_{1}$ in [BG14, equation (9.19)].
@@ 2.4 A formula for the series of planar bipartite maps with tight boundaries
Our input is the Collet-Fusy formula [CF12, Theorem 1.1] for the generating function of planar bipartite maps with $n$ rooted, not necessarily tight, boundaries of prescribed lengths. In our present notations, it reads
@@
for any $n\geq 1$ and positive integers $L_{1},\ldots,L_{n}$. Note that the formula makes sense for $n=1$, upon understanding that $\frac{\partial^{-1}}{\partial t^{-1}}$ means integrating over $t$. For $n=2$, it is a simple exercise to check that the formula is consistent with Corollary 2.3 and (25). We concentrate on the case $n\geq 3$ from now on.
We will deduce from the Collet-Fusy formula an expression for $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\rvert_{\mathrm{bip}}$, using Corollary 2.8. Interestingly, our expression involves a certain family of polynomials that we encountered previously in [BGM24]. Namely, for any integers $k\geq 0$ and $n\geq 1$, let us consider the multivariate polynomial in the variables $\ell_{1},\ldots,\ell_{n}$
@@
where, on the right-hand side, all factors except the first one are $q_{k_{i}}$’s, and where the univariate polynomials $p_{k}$ and $q_{k}$ are defined by
@@
with the convention $p_{0}(\ell)=q_{0}(\ell)=1$, and with $\binom{x}{k}=x(x-1)\cdots(x-k+1)/k!$ viewed as a polynomial in $x$. Clearly, $p_{k}(\ell_{1},\ell_{2},\ldots,\ell_{n})$ is a polynomial in the variables $\ell_{1}^{2},\ldots,\ell_{n}^{2}$. Moreover, as discussed in [BGM24, Proposition 2.1], it is a symmetric polynomial and it satisfies the consistency relation
@@
Combinatorially, $B_{n,k}(r_{1},r_{2},\ldots)$ is the generating series of partitions of an $n$-element set into $k$ blocks, where we attach a weight $r_{i}$ to each block of size $i$. Table 1 lists the first few Bell polynomials. They appear in the following classical formula:
@@ 44 $r_{4}$ $4r_{1}r_{3}+3r_{2}^{2}$ $6r_{1}^{2}r_{2}$ $r_{1}^{4}$
See for instance the discussion in [FS09, Example III.24]. We may now state the main result of this section:
@@
For $n=3$ we recover $T^{(0)}_{2\ell_{1},2\ell_{2},2\ell_{3}}\rvert_{\mathrm{bip}}=R^{\ell_{1}+\ell_{2}+\ell_{3}-1}R^{\prime}-t^{-1}\delta_{\ell_{1}+\ell_{2}+\ell_{3},0}$ consistently with [BGM22, Theorem 1.1], while for $n=4$ we find
@@
When setting the weights $t_{2},t_{4},\ldots$ for inner faces to zero, we have $R=t$, $R^{\prime}=1$ and $R^{(j)}=0$ for $j\geq 2$. Thus, in this case we have $b_{n,k}=t^{-k}\delta_{n,k}$, hence $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}$ reduces to $(n-3)!p_{n-3}(\ell_{1},\ldots,\ell_{n})t^{\ell_{1}+\ldots+\ell_{n}-n+2}$ if at least one $\ell_{i}$ is non-zero, and vanishes otherwise since $p_{n-3}(0,\ldots,0)=(-1)^{n-3}$ so the rightmost term in (33) cancels the result. This is consistent with [BGM24, Theorem 2.3] since the tight maps considered in this reference are nothing but maps with tight boundaries and without inner faces. Note that the exponent of $t$ is consistent with Euler’s relation.
@@
Let us first consider the case where all $\ell_{i}$ are zero: then $T^{(0)}_{0,\ldots,0}$ is simply the generating function of bipartite planar maps with $n\geq 3$ marked vertices. For $n=3$ it is equal to $\frac{\partial}{\partial t}\ln(R/t)$, see e.g. [BGM22, Appendix A], and for larger $n$ we simply have to take more derivatives with respect to $t$, namely
@@
Let us now assume that at least one $\ell_{i}$ is non-zero, and let $m\geq 1$ be the number of non-zero $\ell_{i}$, and s=nms=n-m the number of zero $\ell_{i}$. Without loss of generality, we can assume that $\ell_{1},\ldots,\ell_{m}>0$ and $\ell_{i}=0$ for $i>m$. We note that, in the bipartite case, the trumpet generating function of Proposition 2.4 vanishes whenever its indices are of different parities (since $S=0$), and reads for even indices
@@
where $(L)_{k}:=L(L-1)\cdots(L-k+1)$ denotes the falling factorial. We will rewrite this expression in the same form as the right-hand side of (39), so as to identify $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{m},0,\ldots,0}\right|_{\mathrm{bip}}$. To this end, we proceed as in [BGM24, Section 3.1] and use the Chu-Vandermonde identity to write
@@
Now, by [BGM24, Lemma 3.1] and by (38) [^5] Beware that the notation $A_{\ell,m}$ used in [BGM24] differs from that of the present paper., we have for $L>0$
@@
and by (29), we may replace $p_{k-1}(\ell_{1},\ldots,\ell_{m})$ by $p_{k-1}(\ell_{1},\ldots,\ell_{m},0,\ldots,0)$, in order to have $n$ variables. Comparing this expression with (39), using the invertibility of the unitriangular matrix $(A_{2L,2\ell})_{L,\ell\geq 1}$ and doing a change of variable $k-1\to k$ finally give the wanted formula (33), since the term $\frac{(-1)^{n}(n-3)!}{t^{n-2}}$ is absent for $m\geq 1$. ∎
A variant of the Collet-Fusy formula (26) holds in the quasi-bipartite case where exactly two of the boundaries have odd lengths, i.e. when exactly two of the $L_{i}$, say $L_{1}$ and $L_{2}$, are half-integers. In this case, the factors $\binom{2L_{1}}{L_{1}}\binom{2L_{2}}{L_{2}}$ appearing on the right-hand side should be replaced by $4\binom{2L_{1}-1}{L_{1}-1/2}\binom{2L_{2}-1}{L_{2}-1/2}$. Correspondingly, we have the following analogue of Theorem 2.12:
@@
where, following [BGM24, Section 2.1.3], we define
@@
Starting with the quasi-bipartite variant of the Collet-Fusy formula, we proceed as in the proof of Theorem 2.12. Using Faà di Bruno’s formula, we obtain the expression (40) with the binomial factors $\binom{2L_{1}}{L_{1}}\binom{2L_{2}}{L_{2}}$ modified as discussed above. Now, we expand the falling factorial as follows:
@@
which leads to a variant of (42) in which the factors $\binom{L_{1}-1}{k_{1}}\binom{2L_{1}}{L_{1}}\binom{L_{2}}{k_{2}}\binom{2L_{2}}{L_{2}}$ are replaced by $4\binom{L_{1}-1/2}{k_{1}}\binom{2L_{1}-1}{L_{1}-1/2}\binom{L_{2}-1/2}{k_{2}}\binom{2L_{2}-1}{L_{2}-1/2}$. By [BGM24, Equation (3.9)], we have for $L$ a positive half-integer
@@
Let us consider the case of four boundaries. We have $\tilde{p}_{1}(\ell_{1},\ell_{2};\ell_{3},\ell_{4})=\ell_{1}^{2}+\cdots+\ell_{4}^{2}-\frac{1}{2}$ and $\tilde{p}_{0}(\ell_{1},\ell_{2};\ell_{3},\ell_{4})=1$, hence from Proposition 2.14 we see that, in the quasi-bipartite case, the term 1-1 appearing in the expression (35) should be replaced by $\frac{1}{2}$. In contrast, we found after a lengthy computation (which we omit here) that (35) holds as is when all $\ell_{i}$ are half-integers, i.e. when all four boundaries have odd lengths. This is consistent with the discussion of the lattice count polynomial $N_{0,4}$ at the end of [Nor10, Section 1], since $T^{(0)}_{2\ell_{1},2\ell_{2},2\ell_{3},2\ell_{4}}$ reduces to $N_{0,4}(2\ell_{1},2\ell_{2},2\ell_{3},2\ell_{4})$ when we set the weights for inner faces to zero.